The demand for labor in a certain industry is ND = 300 - w, where ND is the number of workers employers want to hire and w is the real wage measured in dollars per day. The supply of labor in the same industry is NS = 200 + w, where NS is the number of people willing to work. If the minimum wage is set at $60 per day, how many workers will be unemployed?

Select one:

a. 10

b. 20

c. 30

d. 40
Answer #1

At minimum wage of $60 per day.

Demand for labor , ND= 300-w= 300-60= 240 workers

Supply for labor , NS= 200+w = 200+60= 260 workers

Therefore, (260-240)= 20 workers will be unemployed.

Hence, option(B) is correct.